# A probabilistic, data-driven closure model for RANS simulations with aleatoric, model uncertainty

### v1

@@ 1 Introduction
Turbulence is ubiquitous in fluid flows and of importance to a vast range of applications such as aircraft design, climate and ocean modeling. It has challenged and intrigued scientists and artists for centuries Marusic and Broomhall (2021). In the context of the Navier-Stokes equations, the most accurate numerical solution strategy for turbulent flows is offered by Direct Numerical Simulation (DNS), which aims at fully resolving all scales of motion. While this simulation method yields impeccable results, it is prohibitively expensive in terms of computational cost due to the very fine discretizations needed which scale as $\mathcal{O}(\mathrm{Re}^{11/4})$ Pope (2000). Reynolds-averaged Navier-Stokes (RANS) models offer a much more efficient alternative for predicting mean flow quantities. They represent the industry standard which is expected to remain the case in the coming decades Slotnick et al. (2014). Their predictive accuracy however hinges upon the closure model adopted.
@@
Closure models are of three types: (i) Functional, which use physical insight to construct the closure, (ii) Structural, which use mathematical tools, and (iii) Data-driven, which employ experimental/simulation data Ahmed et al. (2021). For a comprehensive review, the reader is directed to San and Maulik (2017); Ahmed et al. (2021); Snyder et al. (2022). The greater availability of computational resources and the development of scalable learning frameworks in the field of machine learning have had a significant impact in computational fluid mechanics as well Brunton et al. (2020); Vinuesa and Brunton (2022); Lucor et al. (2022). Data-driven closures for RANS have revitalized turbulence modeling Duraisamy (2021) and a comprehensive review can be found in Duraisamy et al. (2019); Brunton et al. (2020). The construction of such closure models consist of two steps: (i) postulating a model form ansatz; and (ii) fitting/learning/inferring model parameters on the basis of the available data. Pertinent approaches have focused on learning model coefficients of a given turbulence model Oliver and Moser (2011) (often with statistical inference), on modeling of correction or source terms for an existing turbulence model Parish and Duraisamy (2016); Singh et al. (2017); Tracey et al. (2015); Xiao et al. (2016); Zhang and Duraisamy (2015) and on directly modeling the Reynolds stress (RS) tensor Ling et al. (2016b, a); Kaandorp and Dwight (2020); Wang et al. (2017a) with symbolic regression Schmelzer et al. (2020) or neural networks Ling et al. (2016b); Kaandorp and Dwight (2020); Zhang et al. (2022) or Gaussian Processes Zhang and Duraisamy (2015) or Random Forests Ling et al. (2016a); Wang et al. (2017a). Of particular relevance to the present study is the work of Ling et al. (2016b) wherein they use the non-linear eddy viscosity model (NLEVM) Pope (1975) to capture the anisotropic part of the RS tensor using an integrity tensor basis and a deep neural network employing local, invariant flow features. This model owing to its guaranteed Galilean invariance found a wider utilization Kaandorp and Dwight (2020); Geneva and Zabaras (2019); Zhang et al. (2022). Similarly however to the most widely used RANS closure models, such as the Launder-Sharma $k-\epsilon$ Launder and Sharma (1974) or Wilcox’s $k-\omega$ Wilcox (2008), which are based on the Boussinesq turbulent-viscosity hypothesis, it also postulates that the RS tensor at each point in the problem domain depends on the flow features at the same point (locality assumption). This is a very strong assumption for flows that exhibit strong inhomogeneity Pope (2000).
@@
In most of the methods discussed above, data-based training is performed in an non-intrusive manner, i.e., without involving the RANS solver in the training process. The major shortcomings of such a strategy (which we attempt to address in the present paper) are two-fold. Firstly inconsistency issues, which can arise between the data-driven model and the baseline turbulence model (e.g., $k-\epsilon$) Taghizadeh et al. (2020); Duraisamy (2021). Thompson et al. (2016) showed that even substituting RS fields from reputable DNS databases may not lead to satisfactory prediction of the velocity field, and Wu et al. (2019) investigated the ill-conditioning that arises in the RANS equations, when employing data-driven models that treat the Reynolds stress as an explicit source term. This ill-conditioning can be amplified within each iteration, thus potentially leading to divergence during the solution procedure. Secondly, such models rely on full-field Reynolds stress training data, which are only available when high-fidelity simulations such as DNS/Large-Eddy Simulations (LES) are used. Unfortunately, such high-fidelity simulations due to their expense are limited to simple geometries and low Reynolds numbers.
@@
In order to address these limitations, we advocate incorporating the RANS model in the training process. This enables one to use indirect data (e.g., mean velocities and pressure) obtained from higher-fidelity simulations or experiments as well as direct data (i.e. RS tensor observables) if this is available. In the subsequent discussions, we will refer to such a training strategy as "model-consistent learning" Duraisamy (2021). It necessitates the solution of a high-dimensional inverse problem that minimizes a discrepancy measure between the RANS solver’s output (mean velocities and pressure) and the indirect observations (e.g. mean fields from LES/DNS). Model-consistent or simulation-based Inference Cranmer et al. (2020) further requires that the solver is differentiable in order to provide derivatives of the outputs with respect to the tunable parameters that can in turn be used in the learning/inference algorithm.
@@
In recent years there has been a concerted effort towards developing differentiable CFD solvers Bezgin et al. (2021); List et al. (2022); Um et al. (2021); Kochkov (2021) in Auto-Differentiation (AD) enabled modules like PyTorch, Tensorflow, JAX, Julia. To the best of the authors’ knowledge, this has not been accomplished yet for RANS solvers. One way to enable the computation of parametric sensitivities is by developing adjoint solvers Giles et al. (2003); Giles and Pierce (2000), which are commonly used in the context of aerodynamic shape optimization Jameson (1988). Such adjoint-based modules have also been employed to infer a spatial, corrective field for transport equations Parish and Duraisamy (2015, 2016); Singh et al. (2017) and Reynolds stresses Xiao et al. (2016). Recently, Holland et al. ; Bidar et al. (2022) tried to learn a corrective, multiplicative field in the production term of the Spalart–Allmaras transport model. This is based on an alternative approach outlined in Parish and Duraisamy (2016), in which empirical correction terms for the turbulence transport equations are learned while retaining a traditional linear eddy viscosity model (LEVM) for the closure. Brenner et al. (2022) used adjoints to recover a spatially varying eddy viscosity correction factor from sparsely distributed training data, but they also retained the LEVM assumption. More recently, Ströfer and Xiao (2021) (with adjoint methods) and Zhang et al. (2022) (with ensemble methods) combined the RANS solver and a NLEVM-based neural network proposed by Ling et al. (2016b) in order to learn the underlying model closure. However, they did not account for potential model errors in the closure equation which may arise due to the reasons discussed in the next paragraph.
@@
We argue that even in model-consistent training, a discrepancy in the learnt RS closure model can arise due to the fact that a) the parametric, functional form employed may be insufficient to represent the underlying model [^1] For example the models based on the Boussinesq hypothesis will fail to capture the flow features driven by the anisotropy of the Reynolds stresses and this intrinsic deficiency cannot be remedied by the calibration of the model coefficients with data., and b) the flow features which are used as input in the closure relation and which are generally restricted to each point in the problem domain (locality/Markovianity assumption Parish and Duraisamy (2017)), might not contain enough information to predict the optimal RS tensor leading to irrecoverable loss of information. Hence and irrespective of the type and amount of training data available, there could be aleatoric uncertainty in the closure model that needs to be quantified and propagated in the predictive estimates. We note that very few efforts have been directed towards quantifying uncertainties in RANS turbulence models. Earlier, parametric approaches broadly explored the uncertainties in the model choices ( i.e., uncertainty involved in choosing the best model among a class of competing models, e.g., $k-\omega,k-\epsilon$) and their respective model coefficients. Recently the shortcomings of the parametric closure models have been recognized by the turbulence modeling community Soize (2005). In light of this, various non-parametric approaches have targeted model-form uncertainty whereby uncertainties are directly introduced into the turbulent transport equations or the modeled terms such as the Reynolds stress Geneva and Zabaras (2019) or eddy viscosity. Such formulations allow for more general estimates of the model inadequacy than the parametric approaches. Researchers have also tried perturbing the eigenvalues Emory et al. (2013); Gorlé and Iaccarino (2013); Edeling et al. (2018), transport eigenvectors Thompson et al. (2019) or the tensor invariants. Wu et al. (2017) used kernel density estimates to predict the confidence of data-driven models, but it is limited to the prediction of the anisotropic stress and fails to provide any probabilistic bounds. Geneva and Zabaras (2019) tries to address this issue by incorporating a Bayesian formulation in order to quantify epistemic uncertainty and then propagating it to quantities of interest like pressure and velocity. For a comprehensive review of modeling uncertainties in the RANS models the reader is directed to Xiao and Cinnella (2019).
@@
In order to address the aforementioned limitations, we propose a novel probabilistic, model-consistent, data-driven differential framework. The framework enables learning of a NLEVM-based, RS model in a model-consistent way using a differentiable RANS solver, with mean field observables (velocities and/or pressure). To the authors’ knowledge, uncertainty quantification has not been addressed for data-driven turbulence model training with indirect observations. We propose to augment the parametric closure model of the RS tensor by a stochastic discrepancy tensor to quantify model errors at different parts of the problem domain. With the introduction of the stochastic discrepancy tensor, we advocate a probabilistic formulation for the associated inverse problem, which provides a superior setting as it is capable of quantifying predictive uncertainties which are unavoidable when any sort of model/dimensionality reduction is pursued and when the model (or its closure) is learned from finite data Koutsourelakis et al. (2016). To achieve the desired goals, the proposed framework employs the following major elements:
@@
  • The RS closure model consists of a parametric part that is expressed with an invariant neural network as proposed in Ling et al. (2016b) (Section (2.2.1)), to which a stochastic discrepancy tensor field is added in order to account for the insufficiency of the parametric part (Section (2.2.2)).
  • A fully Bayesian formulation that enables the quantification of epistemic uncertainties and their propagation to the predictive estimates (Section (2.2.4), Section (2.2.5)).
@@
The structure of the rest of the paper is as follows. Section (2) presents the governing equations and their discretization, the closure model proposed consisting of the parametric part and the stochastic corrections provided by latent variables introduced. We also present associated prior and posterior densities, a stochastic Variational Inference scheme that was employed for identifying model parameters and variables as well as the computation of predictive estimates with the trained model. Finally Section (3) discusses the implementation aspects and demonstrates the accuracy and efficacy of the proposed framework in the backward-facing step test case Gresho et al. (1993), where the linear eddy viscosity models are known to fail. We compare our results with LES reference values and the $k-\epsilon$ model which is arguably the most commonly used RANS model. In Section (4), we summarize our findings and discuss limitations and potential enhancements.
@@ 2.1.1 Reynolds-Averaged Navier-Stokes (RANS) equations
The Navier-Stokes equations for incompressible flows of Newtonian fluids are given by:
$\frac{\partial U_{i}}{\partial t}+\frac{\partial}{\partial x_{j}}(U_{i}U_{j})$ $=\nu\frac{\partial^{2}U_{i}}{\partial x_{j}\partial x_{j}}-\frac{1}{\rho}\frac{\partial P}{\partial x_{i}}$ (1)
$\frac{\partial U_{i}}{\partial x_{i}}$ $=0$ (2)
where $U_{i}$, $P$, $t$, $x_{j}$, $\nu$ and $\rho$ represent the flow velocity, pressure, time, spatial coordinates, the dynamic viscosity and the density of the fluid respectively. The non-linearity of the convective term $\frac{\partial}{\partial x_{j}}(U_{i}U_{j})$ gives rise to chaotic solutions beyond a critical value of the Reynolds number $Re$. This necessitates very fine spatio-temporal discretizations in order to capture the salient scales. Such brute-force, fully-resolved simulations, commonly referred to as Direct Numerical Simulations (DNS), can become prohibitively expensive, particularly as $Re$ increases.
@@
Substituting these decompositions into the Navier-Stokes equations (Equation (1)) and applying time-averaging results in the Reynolds-averaged Navier-Stokes (RANS) equations Pope (2000); Alfonsi (2009), i.e.:
@@
$\frac{\partial u_{i}}{\partial x_{i}}$ $=0$ (8)
@@ 2.1.2 The closure problem
$\bm{\tau}_{RS}=-\left<\tilde{u}_{i}\tilde{u}_{j}\right>$ (9)
The goal of pertinent efforts is therefore to devise appropriate closure models where the RS tensor $\bm{\tau}_{RS}$ is expressed as a function as the primary state variables in the RANS equations i.e. the time-averaged flow quantities. Classically, turbulence models are devised to represent higher-order moments of the velocity fluctuations in terms of lower-order moments. This can be done directly, as in the case of the eddy-viscosity models, or indirectly, as in the case of models based on the solution of additional partial differential equations Pope (2000).
@@
$\bm{\tau}_{LEVM}=\frac{2}{3}k\bm{I}-2\nu_{t}\bm{\bar{S}}$ (10)
where $\nu_{t}$ is the eddy viscosity, $\bm{\bar{S}}=\frac{1}{2}\left(\bm{\nabla}\bm{u}+\bm{\nabla}\bm{u}^{T}\right)$ is the mean strain-rate tensor, $\bm{I}$ is the second order identity tensor, and $k=-\frac{1}{2}\text{tr}(\bm{\tau}_{RS})$ is the turbulent kinetic energy. The eddy viscosity is computed after solving the equation(s) for the turbulent flow quantities such as the turbulent kinetic energy $k$ and the turbulent energy dissipation $\epsilon$ (e.g. the $k-\epsilon$ model Launder and Sharma (1974)), or the specific dissipation $\omega$ (e.g. the $k-\omega$ Wilcox (2008)). Although the Boussinesq approximation provides accurate results for a range of flows, it can give rise to predictive inaccuracies which are particularly prominent when trying to capture flows with significant curvatures, recirculation zones, separation, reattachment, anisotropy etc Pope (2000); Wilcox et al. (1998). Attempts to overcome this weakness have been made in the form of nonlinear eddy viscosity models (e.g., Speziale (1987); Pope (1975)) or Reynolds-stress transport models (e.g., Launder et al. (1975)). These models have not received widespread attention because they lack the robustness of LEVM and involve more parameters that need to be calibrated.

### v2

@@ 1 Introduction
Turbulence is ubiquitous in fluid flows and of importance to a vast range of applications such as aircraft design, climate and ocean modeling. It has challenged and intrigued scientists and artists for centuries [1]. In the context of the Navier-Stokes equations, the most accurate numerical solution strategy for turbulent flows is offered by Direct Numerical Simulation (DNS), which aims at fully resolving all scales of motion. While this simulation method yields impeccable results, it is prohibitively expensive in terms of computational cost due to the very fine discretizations needed which scale as $\mathcal{O}(\mathrm{Re}^{11/4})$ [2]. Reynolds-averaged Navier-Stokes (RANS) models offer a much more efficient alternative for predicting mean flow quantities. They represent the industry standard which is expected to remain the case in the coming decades [3]. Their predictive accuracy however hinges upon the closure model adopted.
@@
The greater availability of computational resources and the development of scalable learning frameworks in the field of machine learning have had a significant impact in computational fluid mechanics as well [4, 5, 6]. Data-driven closures for RANS have revitalized turbulence modeling [7], and a comprehensive reviews can be found in [8, 4]. The construction of such closure models consists of two steps: (i) postulating a model form ansatz; and (ii) fitting/learning/inferring model parameters on the basis of the available data. Pertinent approaches have focused on learning model coefficients of a given turbulence model [9] (often with statistical inference), on modeling of correction or source terms for an existing turbulence model [10, 11, 12, 13, 14] and on directly modeling the Reynolds stress (RS) tensor [15, 16, 17, 18] with symbolic regression [19] or neural networks [15, 17, 20] or Gaussian Processes [14] or Random Forests [16, 18]. Of particular relevance to the present study is the work of [15] wherein they use the non-linear eddy viscosity model (NLEVM) [21] to capture the anisotropic part of the RS tensor using an integrity tensor basis and a deep neural network employing local, invariant flow features. This model owing to its guaranteed Galilean invariance found a wider utilization [17, 22, 20].
@@
In most of the methods discussed above, data-based training is performed in a non-intrusive manner, i.e., without involving the RANS solver in the training process. The major shortcomings of such a strategy (which we attempt to address in the present paper) are two-fold. Firstly inconsistency issues, which can arise between the data-driven model and the baseline turbulence model (e.g., $k-\epsilon$) [23, 7]. [24] showed that even substituting RS fields from reputable DNS databases may not lead to satisfactory prediction of the velocity field, and [25] investigated the ill-conditioning that arises in the RANS equations, when employing data-driven models that treat the Reynolds stress as an explicit source term. This ill-conditioning can be amplified within each iteration, thus potentially leading to divergence during the solution procedure. Secondly, such models rely on full-field Reynolds stress training data, which are only available when high-fidelity simulations such as DNS/Large-Eddy Simulations (LES) are used. Unfortunately, such high-fidelity simulations due to their expense are limited to simple geometries and low Reynolds numbers.
@@
In order to address these limitations, we advocate incorporating the RANS model in the training process. This enables one to use indirect data (e.g., mean velocities and pressure) obtained from higher-fidelity simulations or experiments as well as direct data (i.e. RS tensor observables) if this is available. In the subsequent discussions, we will refer to such a training strategy as "model-consistent learning" [7]. It necessitates the solution of a high-dimensional inverse problem that minimizes a discrepancy measure between the RANS solver’s output (mean velocities and pressure) and the observables (e.g. mean fields from LES/DNS). As pointed out in [7], model-consistent training or simulation-based Inference [26] benefits from the differentiability of the solver, as it provides derivatives of the outputs with respect to the tunable parameters that can significantly expedite the learning/inference process.
@@
In recent years there has been a concerted effort towards developing differentiable CFD solvers [27, 28, 29, 30] in Auto-Differentiation (AD) enabled modules like PyTorch, Tensorflow, JAX, Julia. To the best of the authors’ knowledge, this has not been accomplished yet for RANS solvers. One way to enable the computation of parametric sensitivities is by developing adjoint solvers [31, 32], which are commonly used in the context of aerodynamic shape optimization [33]. Such adjoint-based modules have also been employed to infer a spatial, corrective field for transport equations [34, 10, 11] and Reynolds stresses [13]. Recently, [35, 36] tried to learn a corrective, multiplicative field in the production term of the Spalart–Allmaras transport model. This is based on an alternative approach outlined in [10], in which empirical correction terms for the turbulence transport equations are learned while retaining a traditional linear eddy viscosity model (LEVM) for the closure. [37] used adjoints to recover a spatially varying eddy viscosity correction factor from sparsely distributed training data, but they also retained the LEVM assumption. More recently, researchers performed model-consistent or CFD-driven training by involving the solver in the training process. [38, 39] used gradient-free algorithms to perform symbolic identification of the explicit algebraic Reynolds stress models (ARSM) , [40] (with adjoint methods) and [20] (with ensemble methods) combined the RANS solver and a NLEVM-based neural network proposed by [15] in order to learn the model closure. However, they did not account for potential model errors in the closure equation which may arise due to the reasons discussed in the next paragraph.
@@
We argue that even in model-consistent training, a discrepancy in the learnt RS closure model can arise due to the fact that a) the parametric, functional form employed may be insufficient to represent the underlying model [^1] For example the models based on the Boussinesq hypothesis will fail to capture the flow features driven by the anisotropy of the Reynolds stresses and this intrinsic deficiency cannot be remedied by the calibration of the model coefficients with data., and b) the flow features which are used as input in the closure relation and which are generally restricted to each point in the problem domain (locality/Markovianity assumption [41]), might not contain enough information to predict the optimal RS tensor leading to irrecoverable loss of information. Hence irrespective of the type and amount of training data available, there could be aleatoric uncertainty in the closure model that needs to be quantified and propagated in the predictive estimates. We note that much fewer efforts have been directed towards quantifying uncertainties in RANS turbulence models. Earlier, parametric approaches broadly explored the uncertainties in the model choices [9] ( i.e., uncertainty involved in choosing the best model among a class of competing models, e.g., $k-\omega,k-\epsilon$) and their respective model coefficients [42]. Recently the shortcomings of the parametric closure models have been recognized by the turbulence modeling community [43, 44].In light of this, various non-parametric approaches have targeted model-form uncertainty whereby uncertainties are directly introduced into the turbulent transport equations or the modeled terms such as the Reynolds stress [22] or eddy viscosity. Such formulations allow for more general estimates of the model inadequacy than the parametric approaches. Researchers have also tried perturbing the eigenvalues [45, 46, 47], transport eigenvectors [48] or the tensor invariants. [49] used kernel density estimates to predict the confidence of data-driven models, but it is limited to the prediction of the anisotropic stress and fails to provide any probabilistic bounds. [22] tries to address this issue by incorporating a Bayesian formulation in order to quantify epistemic uncertainty and then propagating it to quantities of interest like pressure and velocity. For a comprehensive review of modeling uncertainties in the RANS models the reader is directed to [43].
@@
In order to address the aforementioned limitations, we propose a novel probabilistic, model-consistent, data-driven differential framework. The framework enables learning of a NLEVM-based, RS model in a model-consistent way using a differentiable RANS solver, with mean field observables (velocities and/or pressure). To the authors’ knowledge, uncertainty quantification has not been addressed for data-driven turbulence model training with indirect observations. We propose to augment the parametric closure model of the RS tensor by a stochastic discrepancy tensor to quantify model errors at different parts of the problem domain. With the introduction of the stochastic discrepancy tensor, we advocate a probabilistic formulation for the associated inverse problem, which provides a superior setting as it is capable of quantifying predictive uncertainties which are unavoidable when any sort of model/dimensionality reduction is pursued and when the model (or its closure) is learned from finite data [50]. To achieve the desired goals, the proposed framework employs the following major elements:
@@
  • The RS closure model consists of a parametric part that is expressed with an invariant neural network as proposed in [15] (Section (2.2.1)), to which a stochastic discrepancy tensor field is added in order to account for the insufficiency of the parametric part (Section (2.2.2)).
  • A fully Bayesian formulation that enables the quantification of epistemic uncertainties and their propagation to the predictive estimates (Section (2.2.4), Section (2.2.5))
@@
The structure of the rest of the paper is as follows. Section (2) presents the governing equations and their discretization, the closure model proposed consisting of the parametric part and the stochastic corrections provided by latent variables introduced. We also present associated prior and posterior densities, a stochastic Variational Inference scheme that was employed for identifying model parameters and variables as well as the computation of predictive estimates with the trained model. Finally, Section (3) discusses the implementation aspects and demonstrates the accuracy and efficacy of the proposed framework in the backward-facing step test case [51], where the linear eddy viscosity models are known to fail. We compare our results with LES reference values and the $k-\epsilon$ model, which is arguably the most commonly used RANS model. In Section (4), we summarize our findings and discuss limitations and potential enhancements.
@@ 2.1.1 Reynolds-Averaged Navier-Stokes (RANS) equations
The Navier-Stokes equations for incompressible flows of Newtonian fluids are given by (in indicial notation):
$\frac{\partial U_{i}}{\partial t}+\frac{\partial}{\partial x_{j}}(U_{i}U_{j})$ $=\nu\frac{\partial^{2}U_{i}}{\partial x_{j}\partial x_{j}}-\frac{1}{\rho}\frac{\partial P}{\partial x_{i}},$ (1)
$\frac{\partial U_{j}}{\partial x_{j}}$ $=0,$ (2)
where $i,j$ are free and dummy indices respectively taking values $1,2,3$ and $U_{i}$, $P$, $t$, $x_{j}$, $\nu$ and $\rho$ represent the flow velocity, pressure, time, spatial coordinates, the dynamic viscosity and the density of the fluid respectively. The non-linearity of the convective term $\frac{\partial}{\partial x_{j}}(U_{i}U_{j})$ gives rise to chaotic solutions beyond a critical value of the Reynolds number $Re$. This necessitates very fine spatio-temporal discretizations in order to capture the salient scales. Such brute-force, fully-resolved simulations, commonly referred to as Direct Numerical Simulations (DNS), can become prohibitively expensive, particularly as $Re$ increases.
@@
Substituting these decompositions into the Navier-Stokes equations (Equation (1)) and applying time-averaging results in the Reynolds-averaged Navier-Stokes (RANS) equations [2, 52], i.e.:
@@
$\frac{\partial u_{i}}{\partial x_{i}}$ $=0,$ (8)
@@ 2.1.2 The closure problem
$\bm{\tau}_{RS}=-\left<\tilde{u}_{i}\tilde{u}_{j}\right>.$ (9)
The goal of pertinent efforts is therefore to devise appropriate closure models where the RS tensor $\bm{\tau}_{RS}$ is expressed as a function as the primary state variables in the RANS equations i.e. the time-averaged flow quantities. Closure models are of three types: (i) Functional, which use physical insight to construct the closure; (ii) Structural, which use mathematical tools; and (iii) Data-driven, which employ experimental/simulation data [53]. For a comprehensive review, the reader is directed to [54, 53, 55]. Classically, turbulence models are devised to represent higher-order moments of the velocity fluctuations in terms of lower-order moments. This can be done directly, as in the case of the eddy-viscosity models, or indirectly, as in the case of models based on the solution of additional partial differential equations [2].
@@
$\bm{\tau}_{LEVM}=\frac{2}{3}k\bm{I}-2\nu_{t}\bm{\bar{S}},$ (10)
where $\nu_{t}$ is the eddy viscosity, $\bm{\bar{S}}=\frac{1}{2}\left(\bm{\nabla}\bm{u}+\bm{\nabla}\bm{u}^{T}\right)$ is the mean strain-rate tensor, $\bm{I}$ is the second order identity tensor, and $k=-\frac{1}{2}\text{tr}(\bm{\tau}_{RS})$ is the turbulent kinetic energy. The eddy viscosity is computed after solving the equation(s) for the turbulent flow quantities such as the turbulent kinetic energy $k$ and the turbulent energy dissipation $\epsilon$ (e.g. the $k-\epsilon$ model [56]), or the specific dissipation $\omega$ (e.g. the $k-\omega$ [57]). Although the Boussinesq approximation provides accurate results for a range of flows, it can give rise to predictive inaccuracies which are particularly prominent when trying to capture flows with significant curvatures, recirculation zones, separation, reattachment, anisotropy, etc [2, 58]. Attempts to overcome this weakness have been made in the form of nonlinear eddy viscosity models (e.g., [59, 21, 60]), Reynolds-stress transport models (e.g., [61]) and ARSM (e.g.,[62, 63]). These models have not received widespread attention because they lack the robustness of LEVM and involve more parameters that need to be calibrated.